OpenStudy (anonymous):

If the ratio of a pair of altitudes of similar cones is 3/2, and the volume of the smaller cone is 16 cubic units, what is the volume of the larger cone?

OpenStudy (anonymous):

If the ratios of the linear dimensions of the similar cones is 3/2, then the ratios of their volumes is (3/2)^3 = 27/8 and the volume of the larger cone is (27/8) * 16 = 54 cubic units.